# Most discriminative stimuli for functional cell type identification

### 29 Nov 2023 v1

@@ A.4 Macaque visual area V4 details
Digital twin. The digital twin of macaque V4 neurons, published by Willeke et al. (2023), again followed the core-readout structure. The core of the model consisted of a ResNet50 (He et al., 2016) which was adversarially trained on ImageNet (Deng et al., 2009) to have robust visual representations (Salman et al., 2020). In the digital twin’s core, the first residual block of layer 3 of the ResNet, $layer3.0$, was selected to read out from with $\epsilon=0.1$ because it yielded the highest predictive performance, compared to all other ResNet models and layers. The corresponding size of the output feature map was 1024. The model used batch-normalization (Ioffe & Szegedy, 2015). Lastly, the resulting tensor was rectified with a ReLU unit to obtain the final nonlinear feature space. To predict the response of a single neuron from the feature space, a Gaussian readout was used (Lurz et al., 2020). For each neuron $n$, this readout learned 2d coordinates of the position of the receptive field on the output tensor and extracted a feature vector of length channels at this location. This extracted feature vector was then used in a linear-nonlinear model to predict the neuronal response. To this end, an affine function of the resulting feature vector at the chosen location was computed, followed by a rectifying nonlinearity, chosen to be an ELU (Clevert et al., 2015) offset by one ($\mathrm{ELU}+1$) to make responses positive. The weight vector was $L_{1}$ regularized during training.
@@
Digital twin training. The model was trained to minimize the Poisson loss summed across neurons and computed between observed spike counts and the models’ predicted spike counts with an additional $L_{1}$ regularization of the readout parameters. Models were trained on the full dataset of $n=100$ recording sessions with $n=1244$ neurons and an image size of 100 by 100 pixels. After each epoch, the Poisson loss was computed on the entire validation set. Early stopping was used as follows: If the validation loss did not improve for five consecutive epochs, we restored the weights with the lowest Poisson loss on the validation set and down-scaled the learning rate by a factor of $0.3$ before training continued. After four early stopping steps were completed, the training was stopped.
@@
MDS optimization. We tailored a few MDS optimization parameters to the V4 dataset, optimizing single gray-scale $(100\times 100)$ images with a learning rate of 20, before each optimization step we smoothed image gradients by convolving a Gaussian filter with standard deviation 2 and after each optimization step we normalized the image to a L2-norm of 35. We optimized MDS for 666 steps, and for 1 step during splitting. We stopped MDS optimization early if the exponential moving average of the objective (Eq. 1) did not improve for 100 optimization steps. Furthermore, during splitting we created two new sub-cluster candidates, and terminated splitting after 2 iterations when initializing with 10 clusters and 3 iterations when initializing with 5 clusters (theoretically allowing for the creation of 40 clusters each). As neurons’ responses are not normalized in this data, we post-hoc normalized digital twin predictions such that for each neuron, predictions were standard-normal distributed across the digital twin’s training stimuli, ensuring single cell responses do not dominate the within-cluster mean.
@@ A.5 Extension to two MDS per cluster
Figure 12: Proof of concept of clustering based on two MDS on mouse RGCs. Results on held-out test neurons shown. A. Cluster averaged digital twin response to the optimized MDS. To obtain the response for a cell, the product of the responses to each individual MDS is computed. Elements within a column normalized to highest value. B. Confusion matrix between MDS clusters and Baden et al. (2016) types. Elements within rows (across Baden et al. (2016) types) were normalized to sum to 1, annotated numbers display the number of neurons in percent.
To investigate if our method could be extended to increase cluster granularity by optimizing multiple stimuli, we performed a proof-of-concept experiment: we optimized two MDS per cluster. The responses to both stimuli need to be aggregated before feeding them to the objective (Eq. 1). Here, we chose the product of the responses across stimuli, a soft and differentiable implementation of the “logical and” operation that requires neurons to respond to both stimuli. Initializing with five randomly assigned clusters for the mouse retina, we found a similar result as when optimizing only one stimulus, with a crucial difference: cluster three now showed two different MDS with a sign-flipped temporal UV dependence (Fig. 12A). Only ON-OFF cells would respond to both of these stimuli, and consequently the neurons of this new cluster were mainly associated with ON-OFF Baden et al. (2016) types (Fig. 12B) Interestingly, the MDS also showed orientation-selective features, and hence the according cluster also contained orientation-selective Baden et al. (2016) types.
@@ A.6 Results for another mouse retina digital twin architecture
Figure 13: Most discriminative stimuli clustering on mouse RGCs for another digital twin model architecture. Results on held-out test neurons and empirical comparison to Baden et al. (2016) types shown. A. Cluster averaged digital twin response to the optimized MDS. Elements within a column normalized to highest value. B. Confusion matrix between MDS clusters and Baden et al. (2016) types. Elements within rows (across Baden et al. (2016) types) were normalized to sum to 1, annotated numbers display the number of neurons in percent. C. Confusion matrix between the MDS cluster assignments for the modified digital twin architecture and MDS cluster assignments reported for the original digital twin used for Fig. 2.
@@ A.7 Analysis of temperature hyper-parameter
For lower/higher temperatures, the objective value and the number of clusters found decreased, indicating worse clustering performance. This results directly from the structure of our softmax objective function (Eq. 1). For high temperatures, $\tau\gg\bar{r}_{k}(x_{c}),\ k=1,2,...,c,...$, the inputs to the objective $\bar{r}_{k}(x_{c})/\tau$ are dominated by the temperature, $\tau$. Hence, the objective becomes insensitive to stimulus-dependent inputs $\bar{r}_{k}(x_{c})$, and clustering performance decreases. For small temperatures, $\tau\ll\bar{r}_{k}(x_{c})$, the inputs will be strongly amplified and the response of a single cluster $\bar{r}_{k^{*}}(x_{c})$ will dominate the objective, making it insensitive to the inputs of all other clusters $\bar{r}_{k\neq k^{*}}(x_{c})$. The resulting MDS will then not suppress all clusters. In conclusion, a good temperature balances these two undesirable regimes.

### 14 Mar 2024 v2

@@ A.4 Macaque visual area V4 details
Digital twin. The digital twin of macaque V4 neurons, published by Willeke et al. (2023), again followed the core-readout structure. The core of the model consisted of a ResNet50 (He et al., 2016) which was adversarially trained on ImageNet (Deng et al., 2009) to have robust visual representations (Salman et al., 2020). In the digital twin’s core, the first residual block of layer 3 (layer-3.0) of the ResNet trained with image perturbation constraint $\epsilon=0.1$ was selected to read out from, because it yielded the highest predictive performance, compared to all other ResNet models and layers. The corresponding size of the output feature map was 1024. The model used batch-normalization (Ioffe & Szegedy, 2015). Lastly, the resulting tensor was rectified with a ReLU unit to obtain the final nonlinear feature space. To predict the response of a single neuron from the feature space, a Gaussian readout was used (Lurz et al., 2020). For each neuron $n$, this readout learned spatial coordinates of the position of the receptive field on the output tensor and extracted a feature vector of length channels at this location. This extracted feature vector was then used in a linear-nonlinear model to predict the neuronal response. To this end, an affine function of the resulting feature vector at the chosen location was computed, followed by a rectifying nonlinearity, chosen to be an ELU (Clevert et al., 2015) offset by one ($\mathrm{ELU}(x)+1$) to make responses positive. The weight vector was $L_{1}$-regularized during training.
@@
Digital twin training. The model was trained to minimize the Poisson loss (Eq. 2) summed across neurons and computed between observed spike counts and the models’ predicted spike counts with an additional $L_{1}$ regularization of the readout parameters. Models were trained on the full dataset of $n=100$ recording sessions with $n=1244$ neurons and an image size of 100 by 100 pixels. After each epoch, the Poisson loss was computed on the entire validation set. Early stopping was used as follows: If the validation loss did not improve for five consecutive epochs, we restored the weights with the lowest Poisson loss on the validation set and down-scaled the learning rate by a factor of $0.3$ before training continued. After four early stopping steps were completed, the training was stopped.
@@
MDS optimization. We tailored a few MDS optimization parameters to the V4 dataset, optimizing single gray-scale $(100\times 100)$ pixel images with a learning rate of 20, before each optimization step we smoothed image gradients by convolving a Gaussian filter with standard deviation 2 and after each optimization step we normalized the image to a $L_{2}$-norm of 35. We optimized MDS for 666 steps, and for 1 step during splitting. We stopped MDS optimization early if the exponential moving average of the objective (Eq. 1) did not improve for 100 optimization steps. Furthermore, during splitting we created two new sub-cluster candidates, and terminated splitting after 2 iterations when initializing with 10 clusters and 3 iterations when initializing with 5 clusters (theoretically allowing for the creation of 40 clusters each). As neurons’ responses are not normalized in this data, we post-hoc normalized digital twin predictions such that for each neuron predictions were standard-normal distributed across the digital twin’s training stimuli, ensuring single cell responses do not dominate the within-cluster mean.
@@ A.5 Extension to two MDS per cluster
Figure 12: Proof of concept of clustering based on two MDS on mouse RGCs. Results on held-out test neurons shown. A. Cluster averaged digital twin response to the optimized MDS. To obtain the response for a cell, the product of the responses to each individual MDS is computed. Elements within a column normalized to highest value. B. Confusion matrix between MDS clusters and Baden et al. (2016) types. Elements within rows were normalized to sum to 1, annotated numbers display the number of neurons in percent.
To investigate if our method could be extended to increase cluster granularity by optimizing multiple stimuli, we performed a proof-of-concept experiment: we optimized two MDS per cluster. The responses to both stimuli need to be aggregated before feeding them to the objective (Eq. 1). Here, we chose the product of the responses across stimuli, a soft and differentiable implementation of the “logical and” operation that requires neurons to respond to both stimuli. Initializing with five randomly assigned clusters for the mouse retina, we found a similar result as when optimizing only one stimulus, with a crucial difference: cluster three now showed two different MDS with a sign-flipped temporal UV dependence (Fig. 12A). Only ON-OFF cells would respond to both of these stimuli, and consequently the neurons of this new cluster were mainly associated with ON-OFF Baden et al. (2016) types (Fig. 12B). Interestingly, the MDS also showed orientation-selective features, and hence the according cluster also contained orientation-selective Baden et al. (2016) types.
@@ A.6 Results for another mouse retina digital twin architecture
Figure 13: Most discriminative stimuli clustering on mouse RGCs for another digital twin model architecture. Results on held-out test neurons and empirical comparison to Baden et al. (2016) types shown. A. Cluster averaged digital twin response to the optimized MDS. Elements within a column normalized to highest value. B. Confusion matrix between MDS clusters and Baden et al. (2016) types. Elements within rows were normalized to sum to 1, annotated numbers display the number of neurons in percent. C. Confusion matrix between the MDS cluster assignments for the modified digital twin architecture and MDS cluster assignments reported for the original digital twin used for Fig. 2. Elements within rows were normalized to sum to 1, annotated numbers display the number of neurons in percent.
@@ A.7 Analysis of temperature hyper-parameter
For lower or higher temperatures, the objective value and the number of clusters we found decreased, indicating worse clustering performance. This results directly from the structure of our softmax objective function (Eq. 1). For high temperatures, $\tau\gg\bar{r}_{k}(x_{c}),\ k=1,2,...,c,...$, the inputs to the objective $\bar{r}_{k}(x_{c})/\tau$ are dominated by the temperature, $\tau$. Hence, the objective becomes insensitive to stimulus-dependent inputs $\bar{r}_{k}(x_{c})$, and clustering performance decreases. For small temperatures, $\tau\ll\bar{r}_{k}(x_{c})$, the inputs will be strongly amplified and the response of a single cluster $\bar{r}_{k^{*}}(x_{c})$ will dominate the objective, making it insensitive to the inputs of all other clusters $\bar{r}_{k\neq k^{*}}(x_{c})$. The resulting MDS will then not suppress all clusters. In conclusion, a good temperature balances these two undesirable regimes.